News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 29 January 2010 9.00am EDT

During the last week two observations were impacted on Jan 25 when the ACIS instrument experienced an unexpected Back End Processor (BEP) reboot. The event resulted in a reversion to the pre-launch version of the ACIS flight software. A real-time procedure was run on Jan 25 to restart the ACIS software housekeeping reporting and to dump the contents of the ACIS software patch memory. The anomalous event had also erased the software patches from memory. A second real-time procedure was run on Jan 25 to reload the patches and reboot the ACIS BEP. A third real-time procedure was completed on Jan 26 to reset the ACIS focal-plane control point to -121 degree C as a final clean-up action following the anomaly. The ACIS team has begun an investigation into the anomaly that will include a careful comparison with previous events of this kind.

The schedule of targets for the next week is shown below.

------------------------------------------
Radiation Belts                     Jan 31
SN2009mk               ACIS-S
KTEri                  HRC-S/LETG
Q0957+561              ACIS-S       Feb  1
PG1115+080             ACIS-S
CheshireCatLens        ACIS-S
CygnusOB2 (2 obs)      ACIS-I
MKN421                 ACIS-I/LETG  Feb  2
Cygnus OB2             ACIS-I
SDSSJ113657.68+4113    ACIS-S
2PIGG_m1347            ACIS-I
Radiation Belts
SNRE0519-69.0          ACIS-S       Feb  3
GS0836-429             ACIS-I
2PIGG_m1480 (2 obs)    ACIS-I
2PIGG_m1441            ACIS-I
SDSSJ105755.66+4825    ACIS-S
Cygnus OB2             ACIS-I
M31                    ACIS-I       Feb  4
MKN421                 ACIS-I/LETG
MKN421                 ACIS-I/LETG
Cygnus OB2 (2 obs)     ACIS-I
Radiation Belts                     Feb  5
XMMUJ0236-5225         ACIS-I       Feb  6
KTEri                  HRC-S/LETG
MKN421 (3 obs)         ACIS-I/LETG
SDSSJ163159.59+2437    ACIS-S       Feb  7
XMMUJ0236-5225         ACIS-I
PG1115+080             ACIS-S
SDSSJ155909.62+3501    ACIS-S
Cygnus OB2             ACIS-I
SAXJ1428.6-5422        ACIS-I       Feb  8
2PIGG_m1347            ACIS-I

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports